### v1

@@ 3.3 Path Monotonicity as a Validation Criterion
We return to the defining property of gradient trajectories in smooth Morse theory. Let $f:M\to\mathbb{R}$ be a smooth function on a Riemannian manifold $M$, and let $\gamma:[0,1]\to M$ be a gradient trajectory with $\gamma(0)=p$ and $\gamma(1)=q$. By definition, $\gamma^{\prime}(t)=\nabla f(\gamma(t))$ for all $t$, where $\nabla f$ is the Riemannian gradient. This implies that $f$ varies monotonically along $\gamma$:
$\frac{d}{dt}f(\gamma(t))=\langle\nabla f(\gamma(t)),\gamma^{\prime}(t)\rangle=\|\nabla f(\gamma(t))\|^{2}>0,$ (17)
assuming $\gamma$ does not pass through critical points. The function increases strictly along ascending trajectories and decreases strictly along descending trajectories. This monotonicity is not merely a local property at each point but a global constraint on the entire path.
In the discrete setting, we translate this insight into an operational criterion. For an edge $e=[i,j]$ to participate in gradient flow, we require not only that the function increases from $i$ to $j$ (for ascending flow) but that the increase reflects genuine geometric variation rather than artificial jumping across the manifold. We validate this by examining alternative paths: if the function truly varies smoothly in the region containing $i$ and $j$, then any path connecting them should exhibit consistent monotonic behavior.
Definition 9 (Path Monotonicity).
@@
$\text{cm}(\hat{y};\gamma)=\frac{\sum_{\ell=0}^{k-1}w_{v_{\ell}v_{\ell+1}}\Delta_{v_{\ell}v_{\ell+1}}\hat{y}}{\sum_{\ell=0}^{k-1}w_{v_{\ell}v_{\ell+1}}|\Delta_{v_{\ell}v_{\ell+1}}\hat{y}|}$, (18)
where $\Delta_{v_{\ell}v_{\ell+1}}\hat{y}=\hat{y}(v_{\ell+1})-\hat{y}(v_{\ell})$ and $w_{v_{\ell}v_{\ell+1}}$ are edge weights. For strictly monotone paths, $\text{cm}(\hat{y};\gamma)=1$ for ascending paths and $\text{cm}(\hat{y};\gamma)=-1$ for descending paths.
@@
For an edge $e=[i,j]$ with edge length $\Delta_{e}$ exceeding a threshold $\tau_{q}$ (typically the $q$-quantile of edge lengths for $q\in[0.75,0.90]$), we compute the shortest alternative path $P(i,j)$ that avoids $e$. If no such path exists, the edge is essential for connectivity and we accept it by default. If $P(i,j)$ exists, we evaluate $\text{cm}(\hat{y};P(i,j))$ and compare against a threshold $\theta\in[0.85,1.0]$. For ascending flow, we require $\text{cm}(\hat{y};P(i,j))\geq\theta$; for descending flow, $\text{cm}(\hat{y};P(i,j))\leq-\theta$. The edge participates in gradient flow only if this monotonicity condition holds.
Definition 10 (Validated Edge Set).
@@
$E_{\text{val}}^{\uparrow}=\{[i,j]\in E:\Delta_{e}\leq\tau_{q}\text{ or }\text{cm}(\hat{y};P(i,j))\geq\theta\},$ (19)
@@
and similarly for descending flow with $\text{cm}(\hat{y};P(i,j))\leq-\theta$. We compute gradient trajectories using only edges in $E_{\text{val}}^{\uparrow}$ for ascending basins and $E_{\text{val}}^{\downarrow}$ for descending basins.
@@
This criterion captures the essential geometric property of gradient trajectories while remaining computationally tractable. Computing the shortest alternative path requires breadth-first search after temporarily removing edge $e$, an $O(|E|)$ operation for sparse graphs. Evaluating path monotonicity involves a single pass through the path vertices, adding negligible overhead. The total cost of validating all long edges is $O(L\cdot|E|)$ where $L$ is the number of long edges, typically a small fraction of $|E|$ for well-chosen $q$.
3.4 Theoretical Justification
We provide an informal justification for why path monotonicity validation recovers geometrically faithful gradient flow in the limit of dense sampling. The argument relies on the manifold hypothesis and general position properties of sampled points.
Suppose the data $X_{n}=\{x_{1},\ldots,x_{n}\}$ are sampled from a smooth probability distribution supported on a $d$-dimensional manifold $M$ embedded in $\mathbb{R}^{D}$, and let $f:M\to\mathbb{R}$ be a smooth function with non-degenerate critical points. For sufficiently large $n$ and appropriate choice of $k=k(n)$ growing with $n$ but satisfying $k/n\to 0$, the k-nearest neighbor graph $G_{n}$ approximates the manifold $M$ in the sense that geodesic distances on $G_{n}$ converge to manifold distances on $M$.
Consider an edge $e=[i,j]$ in $G_{n}$ connecting points $x_{i},x_{j}\in M$ with $f(x_{i})<f(x_{j})$. If this edge lies on a gradient trajectory of $f$, then the curve segment connecting $x_{i}$ to $x_{j}$ follows the gradient direction, implying that $f$ increases monotonically between them. For dense sampling, the shortest path $P(i,j)$ in $G_{n}$ (excluding edge $e$) approximates a curve on $M$ close to the direct connection. Since $f$ varies continuously and increases throughout the region, the function values along $P(i,j)$ increase monotonically as well, yielding $\text{cm}(\hat{y};P(i,j))\approx 1$.
Conversely, suppose edge $e$ creates a shortcut across a valley in the function landscape. The edge connects $x_{i}$ and $x_{j}$ that are close in Euclidean distance but separated by a region where $f$ first decreases then increases. The alternative path $P(i,j)$ must traverse this valley, encountering vertices where $\hat{y}$ first decreases from $\hat{y}(x_{i})$ then increases to $\hat{y}(x_{j})$. This oscillation yields $\text{cm}(\hat{y};P(i,j))<\theta$ for reasonable thresholds $\theta$, causing the validation to reject the edge.
The threshold $\theta<1$ provides robustness to sampling noise and discretization effects. In practice, $\theta\in[0.85,0.95]$ accommodates small oscillations due to finite sample size while maintaining the essential monotonicity requirement. The quantile threshold $q$ balances between preserving connectivity and filtering artifacts, with $q\in[0.75,0.90]$ typically appropriate for k-nearest neighbor graphs where most edges are short and only the tail of the length distribution contains problematic edges.
@@ 3.5 Basin Computation with Validated Gradient Flow
We now describe how validated edges integrate into the basin computation algorithm. The construction proceeds through breadth-first exploration from each local extremum, accepting vertices into the basin only when they satisfy the appropriate monotonicity condition along paths from the extremum.
For a local maximum $M$, we construct the descending basin $D(M)$ through the following procedure. Initialize the basin with $D(M)=\{M\}$ and a queue with $M$. Maintain a distance map recording the hop distance from $M$ and a predecessor map tracking paths. While the queue is non-empty, dequeue a vertex $u\in D(M)$ and examine each neighbor $v\notin D(M)$. If $\hat{y}(v)<\hat{y}(u)$ (descending condition), check whether the edge $[u,v]$ requires validation (length $\Delta_{uv}>\tau_{q}$). For short edges, accept $v$ immediately. For long edges, compute the shortest alternative path $P(u,v)\setminus\{u\}$ from $u$ to $v$ excluding the direct edge, and evaluate $\text{cm}(\hat{y};P(u,v))$. If $\text{cm}(\hat{y};P(u,v))\leq-\theta$, the edge is validated and we accept $v$ into the basin, setting its predecessor to $u$ and adding it to the queue. If validation fails, we reject this edge and continue examining other neighbors of $u$.
@@
The algorithm ensures that every vertex in $D(M)$ is connected to $M$ through a path where function values decrease monotonically. Vertices reachable only through invalidated long edges remain outside the basin, even if they would be included under naive gradient flow. This prevents basin jumping while maintaining computational tractability through lazy evaluation: we validate edges only when they arise in the breadth-first search, avoiding exhaustive validation of the entire graph.
@@
Input: Graph $G=(V,E)$, function $\hat{y}:V\to\mathbb{R}$, local maximum $M$, quantile $q$, threshold $\theta$
Output: Basin $D(M)\subseteq V$
 Compute edge length quantile $\tau_{q}=\text{quantile}(\{\Delta_{e}:e\in E\},q)$
@@
while queue not empty do
@@
     $\text{valid}\leftarrow\text{true}$ {Short edge, accept}
@@
     $P\leftarrow\text{ShortestPath}(u,v,G\setminus\{e\})$
     if $P=\emptyset$ then
      $\text{valid}\leftarrow\text{true}$ {Essential for connectivity}
@@
      $\text{valid}\leftarrow\text{cm}(\hat{y};P)\leq-\theta$ {Check monotonicity}

### v2

@@ 3.3 Path Monotonicity as a Validation Criterion
We return to the defining property of gradient trajectories in smooth Morse theory. Let $f:M\to\mathbb{R}$ be a smooth function on a Riemannian manifold $M$, and let $\gamma:[0,1]\to M$ be a gradient flow trajectory with $\gamma(0)=p$ and $\gamma(1)=q$. For ascending flow, $\gamma^{\prime}(t)=+\nabla f(\gamma(t))$; for descending flow, $\gamma^{\prime}(t)=-\nabla f(\gamma(t))$. In both cases, the function varies monotonically along $\gamma$. For ascending trajectories:
$\frac{d}{dt}f(\gamma(t))=\langle\nabla f(\gamma(t)),\gamma^{\prime}(t)\rangle=\|\nabla f(\gamma(t))\|^{2}>0,$
while for descending trajectories, the derivative has the opposite sign. The key property is that $f$ changes monotonically along the entire path, assuming $\gamma$ does not pass through critical points. This monotonicity is not merely a local property at each point but a global constraint on the trajectory.
In the discrete setting, we translate this insight into an operational criterion. For an edge $e=[i,j]$ to participate in gradient flow, we require not only that the function increases from $i$ to $j$ (for ascending flow) but that the increase reflects genuine geometric variation rather than artificial jumping across the manifold. We validate this by examining alternative paths: if the function truly varies smoothly in the region containing $i$ and $j$, then the shortest path connecting them should exhibit consistent monotonic behavior.
@@
$\text{cm}(\hat{y};\gamma)=\frac{\sum_{\ell=0}^{k-1}w_{v_{\ell}v_{\ell+1}}\Delta_{v_{\ell}v_{\ell+1}}\hat{y}}{\sum_{\ell=0}^{k-1}w_{v_{\ell}v_{\ell+1}}|\Delta_{v_{\ell}v_{\ell+1}}\hat{y}|}$,
where $\Delta_{v_{\ell}v_{\ell+1}}\hat{y}=\hat{y}(v_{\ell+1})-\hat{y}(v_{\ell})$ and $w_{v_{\ell}v_{\ell+1}}$ are edge weights. $\text{cm}(\hat{y};\gamma)=1$ for $\hat{y}$ strictly ascending along $\gamma$ and $\text{cm}(\hat{y};\gamma)=-1$ for $\hat{y}$ strictly descending along $\gamma$.
@@
For an edge $e=[i,j]$ with edge length $\Delta_{e}$ exceeding a threshold $\tau_{q}$ (typically the $q$-quantile of edge lengths for $q\in[0.75,0.90]$), we compute the shortest path $\gamma_{[i,j]}$ from $i$ to $j$ in the graph $G\setminus\{e\}$ (after temporarily removing edge $e$). If no such path exists, the edge is essential for connectivity and we accept it by default. If $\gamma_{[i,j]}$ exists, we evaluate $\text{cm}(\hat{y};\gamma_{[i,j]})$ and compare against a threshold $\theta\in[0.85,1.0]$. For ascending flow, we require $\text{cm}(\hat{y};\gamma_{[i,j]})\geq\theta$; for descending flow, $\text{cm}(\hat{y};\gamma_{[i,j]})\leq-\theta$. The edge participates in gradient flow only if this monotonicity condition holds.
@@
$E_{\text{val}}^{\uparrow}=\{[i,j]\in E:\hat{y}(j)>\hat{y}(i)\text{ and }(\Delta_{e}\leq\tau_{q}\text{ or }\mathcal{V}^{\uparrow}(e))\},$
where $\mathcal{V}^{\uparrow}(e)$ holds if either no path from $i$ to $j$ exists in $G\setminus\{e\}$, or the shortest such path $\gamma_{[i,j]}$ satisfies $\text{cm}(\hat{y};\gamma_{[i,j]})\geq\theta$. Similarly, for descending flow:
$E_{\text{val}}^{\downarrow}=\{[i,j]\in E:\hat{y}(j)<\hat{y}(i)\text{ and }(\Delta_{e}\leq\tau_{q}\text{ or }\mathcal{V}^{\downarrow}(e))\},$
@@
where $\mathcal{V}^{\downarrow}(e)$ holds if either no path from $i$ to $j$ exists in $G\setminus\{e\}$, or $\text{cm}(\hat{y};\gamma_{[i,j]})\leq-\theta$. We compute gradient trajectories using only edges in $E_{\text{val}}^{\uparrow}$ for ascending basins and $E_{\text{val}}^{\downarrow}$ for descending basins.
@@
This criterion captures the essential geometric property of gradient trajectories while remaining computationally tractable. Computing the shortest path in $G\setminus\{e\}$ requires breadth-first search, an $O(|E|)$ operation for sparse graphs. Evaluating path monotonicity involves a single pass through the path vertices, adding negligible overhead. The total cost of validating all long edges is $O(L\cdot|E|)$ where $L$ is the number of long edges, typically a small fraction of $|E|$ for well-chosen $q$.
3.4 Local Extrema and Spurious Feature Removal
We return to a fundamental question that arises whenever discrete Morse theory is applied to finite samples from continuous manifolds: how should we distinguish genuine features of the underlying geometry from artifacts introduced by discretization and statistical estimation? The gradient flow structure computed from path-validated edges naturally partitions the graph into basins of attraction, with each basin containing vertices that flow to a common local extremum. However, not all local extrema identified by this purely combinatorial criterion represent meaningful features of the conditional expectation function. Some extrema arise from sampling variability, others from numerical artifacts in the smoothing process, and still others from the discretization itself, where the restriction of a smooth function to graph vertices inevitably creates spurious critical points that do not correspond to extrema of the continuous function.
Consider first the geometric origin of spurious extrema. Let $f:M\to\mathbb{R}$ be a smooth function on a Riemannian manifold $M$, and let $X=\{x_{1},\ldots,x_{n}\}$ be a finite sample from $M$. The restriction $y=f|_{X}$ defines a function on the sample points, which we represent as a function on the vertices of the graph $G_{k}(X)$ derived from the k-nearest neighbor construction. Even when $f$ has only a small number of critical points on $M$, the discrete function $y$ typically exhibits numerous local extrema on $G_{k}(X)$. This phenomenon occurs because discrete extrema are defined purely by local comparisons: a vertex $v$ is a local minimum if $y(v)<y(u)$ for all neighbors $u\in N(v)$, and a local maximum if $y(v)>y(u)$ for all $u\in N(v)$. The graph structure imposes a specific neighborhood system that may not align with the natural neighborhoods in the continuous manifold, leading to vertices that satisfy the discrete extremum condition despite lying on smooth portions of the function landscape.
After spectral smoothing to estimate the conditional expectation, additional spurious extrema can emerge from the interplay between the smoothing operator and the graph geometry. The heat kernel filter $\exp(-tL_{\text{norm}})$ diffuses function values across edges, with diffusion rates determined by the Riemannian structure. In regions where the graph locally approximates the manifold geometry well, smoothing faithfully reconstructs the underlying smooth function. However, in regions where long edges bridge across valleys or where local density variations create anomalous mass distributions, the smoothed function may develop local extrema that reflect these geometric artifacts rather than true features of the conditional expectation.
We address this challenge through a systematic refinement pipeline that progressively removes spurious extrema based on multiple geometric and statistical criteria. The refinement process operates in stages, each targeting a different source of spuriousness while preserving the structure of the function landscape.
The first filtering stage removes extrema whose function values are insufficiently distinct from the global mean. We compute the relative value $r(v)=\hat{y}(v)/\bar{y}$ for each extremum vertex $v$, where $\bar{y}=n^{-1}\sum_{i=1}^{n}y(i)$ denotes the mean of $y$ across all vertices. For local maxima, we retain only those satisfying $r(v)\geq\rho_{\max}$ for a threshold $\rho_{\max}>1$ (typically $\rho_{\max}\in[1.1,1.5]$), while for local minima we require $r(v)\leq\rho_{\min}$ with $\rho_{\min}<1$ (typically $\rho_{\min}\in[0.5,0.9]$). These thresholds focus subsequent analysis on prominent features that rise substantially above or descend significantly below the average function level. Extrema failing these criteria represent minor fluctuations that, while formally satisfying the local comparison condition, lack sufficient magnitude to warrant interpretation as distinct features of the conditional expectation surface.
The second stage addresses redundancy arising from multiple nearby extrema representing the same underlying feature. We quantify similarity between extrema through their basin overlap. For two basins $A_{i}$ and $A_{j}$ with vertex sets $V_{i}$ and $V_{j}$, the overlap coefficient is defined as
$\omega(A_{i},A_{j})=\frac{|V_{i}\cap V_{j}|}{\min(|V_{i}|,|V_{j}|)},$ (6)
which measures what fraction of the smaller basin’s vertices are shared with the larger basin. This asymmetric measure emphasizes cases where one basin is substantially contained in another, indicating that the corresponding extrema likely represent the same feature at different scales of resolution. We construct an overlap graph where vertices correspond to extrema of the same type (all maxima or all minima), and we add an edge between extrema $i$ and $j$ whenever $\omega(A_{i},A_{j})\geq\omega_{\text{thld}}$ for a threshold $\omega_{\text{thld}}\in[0.10,0.20]$. Connected components in this overlap graph identify clusters of similar extrema. Within each cluster, we merge the basins by retaining only the extremum with the most extreme function value (highest for maxima, lowest for minima) and assigning all vertices in the cluster’s combined basin to this representative extremum. This consolidation reduces redundancy while preserving the essential gradient flow structure, as vertices in merged basins still flow toward genuine extrema, just through a simplified representative structure.
The third stage removes extrema whose basins exhibit geometric characteristics suggesting isolation artifacts or boundary effects. We compute two complementary measures of basin geometry for each extremum. The first is the mean hop-k distance, which quantifies extended neighborhood isolation. For an extremum at vertex $v$, we identify all vertices $U_{k}(v)$ at graph distance exactly $k$ from $v$ through breadth-first search, then compute the mean geodesic distance
$d_{k}(v)=\frac{1}{|U_{k}(v)|}\sum_{u\in U_{k}(v)}d_{G}(v,u),$ (7)
where $d_{G}(v,u)$ denotes the length of the shortest path from $v$ to $u$ in the metric graph. This measure captures whether the extremum lies in a locally sparse or dense region at scale $k$ (typically $k=2$ or $k=3$). High values of $d_{k}(v)$ indicate that vertices at hop distance $k$ are geometrically far from $v$, suggesting the extremum sits in an isolated or boundary region.
The second geometric measure is the effective degree $\deg_{\text{eff}}(v)$, defined as the sum of density-surrogate weights over all neighbors:
$\deg_{\text{eff}}(v)=\sum_{u\in N(v)}w(e_{vu}),$ (8)
where $w(e_{vu})$ denotes the edge weight (often taken as the edge mass $m_{1}([v,u])$). This weighted degree quantifies how well-connected the vertex is to the graph structure, accounting for local density variations. Vertices with anomalously low effective degree relative to the global distribution may represent poorly sampled boundary regions where the discrete gradient flow is unreliable.
We convert both geometric measures to percentile ranks across all vertices: for $d_{k}$, we compute
$p_{k}(v)=|\{u:d_{k}(u)\leq d_{k}(v)\}|/n,$
which gives the fraction of vertices with hop-k distance at most that of $v$. Similarly, for effective degree, we compute
$p_{\deg}(v)=|\{u:\deg_{\text{eff}}(u)\geq\deg_{\text{eff}}(v)\}|/n,$
giving the fraction of vertices with effective degree at least that of $v$ (note the reversed inequality, as high degree indicates good connectivity). We retain extrema only if both geometric measures fall within acceptable ranges: typically $p_{k}(v)<0.90$ and $p_{\deg}(v)>0.10$, though these thresholds may be adjusted based on the specific graph geometry and application requirements.
The complete refinement pipeline applies these stages sequentially: relative value filtering removes extrema with insufficient magnitude, overlap-based clustering and merging consolidates redundant features, and geometric filtering removes isolated or poorly-connected extrema. Each stage preserves the basin structure for retained extrema, so the final refined basin structure maintains complete information about gradient flow for all vertices assigned to surviving extrema. This multi-stage approach balances statistical and geometric considerations, ensuring that the final set of extrema represents genuine features of the estimated conditional expectation function rather than artifacts of discretization, sampling, or smoothing.
The filtering criteria involve several tunable parameters: the relative value thresholds $\rho_{\max}$ and $\rho_{\min}$, the overlap threshold $\omega_{\text{thld}}$, and the geometric percentile thresholds for $p_{k}$ and $p_{\deg}$. In practice, these parameters can be selected through exploratory analysis, examining how the number and prominence of retained extrema vary with parameter choices, or through cross-validation by assessing the stability of the resulting basin structure under perturbations of the smoothed function. Conservative choices (stricter thresholds) produce fewer extrema with higher confidence, while liberal choices preserve more potential features at the cost of retaining some spurious extrema. The modular design of the pipeline allows each filtering stage to be enabled or disabled independently, providing flexibility for different application contexts and data characteristics.
@@ 3.5 Basin Computation with Validated Gradient Flow
We turn now to the computational construction of basins of attraction using validated gradient flow. The basin structure provides the foundation for regional statistical inference, partitioning the sample space into regions where outcomes flow toward common extrema. Before developing the discrete algorithm, we recall the classical construction from smooth Morse theory, which both motivates our approach and highlights a fundamental difference between the continuous and discrete settings that has important implications for statistical applications.
Classical Basins and Gradient Flow Trajectories
In the smooth setting, consider a Morse function $f:M\to\mathbb{R}$ on a compact Riemannian manifold $M$ without boundary. The gradient flow generates two types of trajectories from any regular point $x\in M$ (a point where $\nabla f(x)\neq 0$). The ascending trajectory satisfies the differential equation $\gamma^{\prime}(t)=+\nabla f(\gamma(t))$ with initial condition $\gamma(0)=x$, following the direction of steepest increase until reaching a local maximum as $t\to\infty$. The descending trajectory satisfies $\gamma^{\prime}(t)=-\nabla f(\gamma(t))$, flowing downhill to a local minimum. The Morse-Smale condition (generic transversality of stable and unstable manifolds) ensures that these trajectories are well-defined and unique for almost all starting points, with trajectories terminating at critical points rather than wandering indefinitely.
These trajectories induce maps $\pi_{\uparrow}:M\to\text{Max}(f)$ and $\pi_{\downarrow}:M\to\text{Min}(f)$ that assign to each point the terminus of its ascending and descending flows. For a local minimum $m$, the ascending basin (or unstable manifold) is defined as
$A(m)=\{x\in M:\pi_{\downarrow}(x)=m\},$ (9)
containing all points whose descending flow terminates at $m$. For a local maximum $M$, the descending basin (or stable manifold) is
$D(M)=\{x\in M:\pi_{\uparrow}(x)=M\},$ (10)
containing all points whose ascending flow terminates at $M$. The terminology reflects the direction of flow reaching the critical point: vertices in $A(m)$ descend to reach $m$, while vertices in $D(M)$ ascend to reach $M$.
A crucial property of the smooth setting is that basins of the same type are disjoint. No point can have descending flow terminating at two distinct minima, nor can ascending flow from a single point reach two distinct maxima. This uniqueness follows from the smooth dependence of gradient trajectories on initial conditions and the Morse-Smale transversality assumption. The ascending basins $\{A(m):m\in\text{Min}(f)\}$ partition the manifold into disjoint regions, as do the descending basins $\{D(M):M\in\text{Max}(f)\}$. The intersections $C(m,M)=A(m)\cap D(M)$ define gradient flow cells, which tile the manifold into regions where both ascending and descending flows have unique termini.
Non-Uniqueness in the Discrete Setting
When we discretize the gradient flow by restricting to a finite graph $G=(V,E)$, this uniqueness property fails. The discrete gradient flow is defined by following edges of steepest ascent or descent, but when multiple neighbors have identical or nearly identical function values, the choice of which edge to follow becomes ambiguous. More fundamentally, the graph topology itself creates situations where multiple distinct paths exist between a vertex and an extremum, each exhibiting monotonic function behavior, yet leading to different flow trajectories depending on the algorithmic tie-breaking rules.
Consider a simple example that illustrates this phenomenon. Let $G$ be a star graph with a central vertex $c$ and three arms, each consisting of a single edge connecting $c$ to a terminal vertex. Label the terminal vertices $v_{1}$, $v_{2}$, and $v_{3}$. Define a function $\hat{y}$ with values $\hat{y}(v_{1})=0$, $\hat{y}(v_{2})=0$, $\hat{y}(c)=1$, and $\hat{y}(v_{3})=2$. The vertices $v_{1}$ and $v_{2}$ are both local minima (each has function value less than its only neighbor $c$), while $v_{3}$ is the unique local maximum.
Now consider the ascending basins. The vertex $c$ has function value intermediate between the two minima and the maximum. From $c$, descending flow could reasonably proceed to either $v_{1}$ or $v_{2}$, as both are neighbors with lower function values. If we break ties by choosing the neighbor with minimal function value, both $v_{1}$ and $v_{2}$ qualify equally. Different tie-breaking rules lead to different basin assignments: one rule might assign $c$ to $A(v_{1})$, another to $A(v_{2})$. Regardless of the choice, the resulting ascending basins are not disjoint. If $c\in A(v_{1})$, then since $v_{2}$ is a minimum, we have $A(v_{2})=\{v_{2}\}$. But if we run the basin construction algorithm symmetrically from both minima using breadth-first exploration (as we describe below), the vertex $c$ appears reachable from both $v_{1}$ and $v_{2}$ through monotone descending paths ($v_{1}\to c$ and $v_{2}\to c$ both ascend, so tracing backward from $c$ can descend to either minimum). Thus $c$ potentially belongs to both ascending basins, creating an intersection $A(v_{1})\cap A(v_{2})\ni c$.
This non-uniqueness has important statistical implications. In the continuous setting, the partition into gradient flow cells provides an unambiguous spatial decomposition for regional inference. In the discrete setting, ambiguous vertices near basin boundaries require explicit resolution through tie-breaking rules or probabilistic assignment. Our algorithm addresses this by using breadth-first exploration to establish basin membership definitively based on discovery order, but we acknowledge that alternative algorithmic choices could produce different basin structures for ambiguous vertices. This sensitivity to algorithmic details suggests that statistical inference should focus on vertices deep within basins, where flow direction is unambiguous, rather than on vertices near boundaries where multiple extrema exert comparable influence.
Backward Tracing Through Breadth-First Exploration
We construct basins through breadth-first exploration that works backward along flow trajectories. The key insight is that to find all vertices whose ascending flow reaches a maximum $M$, we start at $M$ itself and explore outward, accepting vertices that lie ”below” already-accepted vertices in function value. This backward tracing ensures that every accepted vertex has a monotone increasing path to $M$, which corresponds precisely to the forward ascending trajectory from that vertex terminating at $M$.
Consider the descending basin $D(M)=\{v:\pi_{\uparrow}(v)=M\}$ of a local maximum $M$. We initialize the basin with $D(M)=\{M\}$ and a queue containing $M$. We maintain a distance map $\text{dist}[v]$ recording the hop distance from $M$ to each accepted vertex, and a predecessor map $\text{pred}[v]$ tracking the path back to $M$. The algorithm proceeds iteratively: while the queue is non-empty, we dequeue a vertex $u\in D(M)$ and examine each neighbor $v\notin D(M)$. If $\hat{y}(v)<\hat{y}(u)$, then the edge $[v,u]$ represents a valid step in an ascending trajectory from $v$ toward $M$ (since moving from $v$ to $u$ increases the function value). To verify that this edge participates in genuine gradient flow rather than an artifactual jump across basins, we apply the validation criterion from Section 3.3.
For edges with length $\Delta_{uv}\leq\tau_{q}$ (below the quantile threshold), we accept $v$ immediately, as short edges are presumed to follow local manifold geometry faithfully. For long edges with $\Delta_{uv}>\tau_{q}$, we compute the shortest alternative path $\gamma_{[u,v]}$ from $u$ to $v$ in the graph $G\setminus\{[u,v]\}$ (temporarily removing the direct edge). If no such path exists, the edge is essential for connectivity and we accept it by default. If an alternative path exists, we evaluate its co-monotonicity coefficient $\text{cm}(\hat{y};\gamma_{[u,v]})$. For descending exploration (where we seek vertices with $\hat{y}(v)<\hat{y}(u)$ that can ascend to $M$), the path from $u$ to $v$ should exhibit descending monotonicity (function values decreasing along the path from $u$ toward $v$). We require $\text{cm}(\hat{y};\gamma_{[u,v]})\leq-\theta$ for a threshold $\theta\in[0.85,1.0]$. If this condition holds, the long edge is validated and we accept $v$ into the basin, setting $\text{dist}[v]=\text{dist}[u]+1$, $\text{pred}[v]=u$, and adding $v$ to the queue for further exploration. If validation fails, we reject this particular edge from $u$ to $v$ and continue examining other neighbors.
@@
This backward exploration guarantees that every vertex in $D(M)$ is connected to $M$ through a validated path where function values increase monotonically. Vertices reachable from $M$ only through invalidated long edges remain outside the basin, preventing spurious basin jumping while maintaining computational efficiency. The validation occurs lazily during breadth-first search, so we evaluate co-monotonicity only for long edges that actually arise in potential basin membership, avoiding exhaustive validation of the entire graph.
@@
Input: Graph $G=(V,E)$, function $\hat{y}:V\to\mathbb{R}$, local maximum $M$, quantile threshold $q$, monotonicity threshold $\theta$
Output: Descending basin $D(M)=\{v:\pi_{\uparrow}(v)=M\}\subseteq V$
 Compute edge length quantile $\tau_{q}\leftarrow\text{quantile}(\{\Delta_{e}:e\in E\},q)$
@@
while queue is not empty do
@@
     $\text{valid}\leftarrow\text{true}$ {Short edge: $v$ can ascend through $u$ to MM}
@@
     $\gamma_{[u,v]}\leftarrow\text{ShortestPath}(u,v,G\setminus\{e\})$
     if $\gamma_{[u,v]}=\emptyset$ then
      $\text{valid}\leftarrow\text{true}$ {Edge essential for connectivity}
@@
      $\text{valid}\leftarrow(\text{cm}(\hat{y};\gamma_{[u,v]})\leq-\theta)$ {Validate descending path monotonicity}